09:44
Run 26 started - Beam no target
Beam energy is being increased from 1.5 to 3 MeV
Horizontal beam position = 7.0 mm, vertical beam position = 0.0 mm.
Detectors are positioned fully OUT
Histograms and stats zero'd, monitoring code restarted.
Pressure in YR09 E target ~ 4.5e-11 mbar.
Merger OK, Tape server OK
ADC data item statistics looks as before. OK
FEE temperatures look OK
System checks OK
Leakage current stable
10:48 (Attachements 1-4)
Merger OK, Tape server OK
ADC data item statistics looks as before. OK
FEE temperatures look OK
System checks OK
Leakage current stable
The detectors currently fully out for tuning.
11:57 (attachments 5-8)
Merger OK, Tape server OK
ADC data item statistics looks as before. OK
FEE temperatures look OK
System checks OK
Leakage current stable
The detectors currently fully out for tuning.
|